OpenStudy (anonymous):

@Michele_Laino A spinner has five sections, labeled A, B, C, D, and E. The spinner is spun 45 times, and the results are recorded in the table. What is the experimental probability of the spinner’s landing on C? Round to the nearest percent, if necessary. Outcome A B C D E Number of trials 6 10 8 12 9 A. 8% B. 18% C. 20% D. 24%

OpenStudy (michele_laino):

hint: \[probability = \frac{{favorable\;cases}}{{possible\;cases}} = \frac{8}{{\left( {6 + 10 + 8 + 12 + 9} \right)}} = ...?\]

OpenStudy (michele_laino):

after that computation, you have to multiply your result by 100. What do you get?

OpenStudy (anonymous):

B?

OpenStudy (michele_laino):

that's right!